JoelTheBoss (joel_the_boss):

What is the equation in point-slope form of the line passing through (4, 0) and (2, 6)?

JoelTheBoss (joel_the_boss):

y = 4x − 2 y = 2x − 4 y = −3(x − 4) y = 3(x + 4)

OpenStudy (marziajackson$):

I got the second answer